# Recurrence method in Non-Hermitian Systems

### v1

@@ VI Conclusion
We have proposed a novel recurrence method for the energy spectra of non-Hermitian systems under open boundary conditions. Based on the recurrence relations of their characteristic polynomials, our formalism transforms the eigenvalue problems of the target multi-band non-Hermitian Hamiltonians to the solutions of a series of polynomials. Therefore, it resolves the accuracy and instability issues in the numerical diagonalization of non-Hermitian matrices, especially for extensive system sizes $N$; with the fast-converging Durand-Kerner method, it also displays better computational complexity than the non-Bloch band theory and the GBZ method, especially for non-Hermitian systems with a considerable period or unit cell with many sublattices. Indeed, as we have demonstrated, it yields consistent analytical spectra or expressions for the non-Hermitian Su-Schrieffer-Heeger models and the non-Hermitian Rice-Mele models under open boundary conditions and efficient and precise numerical results of the non-Hermitian Hofstadter butterfly with a hefty $q_{\text{max}}$ and thus a large magnetic unit cell beyond previous approaches.
The recurrence method also provides a targeted formulation of the non-Hermitian edge spectra, with algorithmic and cost advantages over the previous edge matrix method. We have shown its applications on the edge spectra of multi-band non-Hermitian models with nearest-neighbor hopping and open boundary conditions. For instance, we have studied topological edge modes’ presence and stability (against perturbations) in the non-Hermitian 1D Rice-Mele and the 2D non-Hermitian Hofstadter models, consistent with and convenient for their topological characterizations. In particular, we have also generalized the size-parity effect of the Hermitian Rice-Mele models to non-Hermitian counterparts and without the generalized chiral symmetry.
The interplay between open boundaries and non-Hermitian, non-reciprocal physics has gathered much recent interest and attention and witnessed the discoveries of a series of intriguing and exotic phenomena; the introductions of extra factors such as topological phenomena, magnetic fields, dimensionality, etc., has further developed and enriched these research areas; therefore, our recurrence method offers a timely and promising formalism to facilitate and streamline studies on such non-Hermitian systems under open boundary conditions.
@@ A.1 Properties of Recurrence Relations
Let $M=[h_{ab}]$ be the Hamiltonian matrix of model $H=\sum\limits_{ab}c_{a}^{\dagger}h_{ab}c_{b}$ with hopping range $(l,r)$, namely $h_{ab}\neq 0$ only for $-r\leq a-b\leq l$. Denote the determinant of submatrix containing first $n$ lattice sites as $M_{n}=\det[h_{ab}]_{1\leq a,b\leq n}$. Then for large $n$ $M_{n}$ satisfies a recurrence relation $M_{n}=\sum\limits_{j=1}^{s}R_{j}(\{h_{ab}\})M_{n-j}$ where $s=(l+r)!/l!r!$ is the order of recurrence relation and coefficients $R_{j}$ satisfy $R_{j}(\{xh_{ab}\})=x^{j}R_{j}(\{h_{ab}\})$.
@@
One can use the following algorithm to obtain the recurrence relation coefficients $R$ based on Theorem 1:
@@
The OBC spectrum of a non-Hermitian Hamiltonian $H=\sum\limits_{ij}c_{i}^{\dagger}h_{ij}c_{j}$ is determined from solutions of determinant $\det(E-H)$. Denote the determinant of $N$-site lattice model as $D_{N}:=\det(EI_{N}-H_{N})$, then $D_{N}(E)$ is a polynomial of degree $N$ and satisfies the following recurrence relation according to Theorem 1:
@@ Appendix B Recurrence method for one-dimensional periodic nearest-neighbor hopping model
Note that both $\text{Tr}(U)$ and $\det(U)$ are size-independent, so we can apply the recurrence relation method in the main text to obtain the OBC spectrum. For $k=3$ case with identical nearest-neighbor hopping $t_{1}=t_{2}=t_{3}\equiv u_{1}$, $t_{1}^{\prime}=t_{2}^{\prime}=t_{3}^{\prime}\equiv u_{2}$, the coefficients of the recurrence relation in Eq. (69) are: $\text{Tr}(U)=E^{3}-E^{2}(V_{1}+V_{2}+V_{3})+E(-3u_{1}u_{2}+V_{1}V_{2}+V_{1}V_{3}+V_{2}V_{3})+u_{1}u_{2}(V_{1}+V_{2}+V_{3})-V_{1}V_{2}V_{3}$, $\det(U)=u_{1}^{3}u_{2}^{3}$. For a general $k$, we can show that the recurrence relation is given by:
@@
where
@@ Appendix D Non-Hermitian Hofstadter butterfly on the triangular lattice
where the magnetic phase factors along the three sides of a triangular unit cell are given by

### v2

@@ VI Conclusion
We have proposed a novel recurrence method for the energy spectra of non-Hermitian systems under open boundary conditions. Based on the recurrence relations of their characteristic polynomials, our formalism transforms the eigenvalue problems of the target multi-band non-Hermitian Hamiltonians to the solutions of a series of polynomials. It is generally applicable to non-Hermitian systems with (partial) lattice translation symmetries. Therefore, it resolves the accuracy and instability issues in the numerical diagonalization of non-Hermitian matrices, especially for extensive system sizes $N$; with the fast-converging Durand-Kerner method, it also displays better computational complexity than the non-Bloch band theory and the GBZ method, especially for non-Hermitian systems with a considerable period or unit cell with many sublattices. Indeed, as we have demonstrated, it yields consistent analytical spectra or expressions for the non-Hermitian Su-Schrieffer-Heeger models and the non-Hermitian Rice-Mele models under open boundary conditions and efficient and precise numerical results of the non-Hermitian Hofstadter butterfly with a hefty $q_{\text{max}}$ and thus a large magnetic unit cell beyond previous approaches. These spectra host rich physics regarding the PT-symmetry breaking and topological phenomena of the target non-Hermitian systems.
The recurrence method also provides a targeted formulation of the non-Hermitian edge spectra, with algorithmic and cost advantages over the previous edge matrix method. We have shown its applications on the edge spectra of multi-band non-Hermitian models with nearest-neighbor hopping and open boundary conditions. For instance, we have studied topological edge modes’ presence and stability (against perturbations) in the non-Hermitian 1D Rice-Mele and the 2D non-Hermitian Hofstadter models, consistent with and convenient for their topological characterizations - also as examples of the bulk-boundary correspondences. In particular, we have also generalized the size-parity effect of the Hermitian Rice-Mele models to non-Hermitian counterparts and without the generalized chiral symmetry.
The interplay between open boundaries and non-Hermitian, non-reciprocal physics has gathered much recent interest and attention and witnessed the discoveries of a series of intriguing and exotic phenomena; the introductions of extra factors such as topological phenomena, magnetic fields, dimensionality, etc., has further developed and enriched these research areas; therefore, our general, efficient, and accurate recurrence method offers a timely and promising formalism to facilitate and streamline studies on such non-Hermitian systems under open boundary conditions. The codes and results of this paper are available at Github  [^2] For codes and corresponding results in the Hofstadter model see https://github.com/chenhaoyan05/recurrence_method.
@@ A.1 Properties of Recurrence Relations
Let $M=[h_{ab}]$ be the Hamiltonian matrix of model $H=\sum\limits_{ab}c_{a}^{\dagger}h_{ab}c_{b}$ with hopping range $(l,r)$, namely $h_{ab}\neq 0$ only for $-r\leq a-b\leq l$. Denote the determinant of submatrix containing first $n$ lattice sites as $M_{n}=\det[h_{ab}]_{1\leq a,b\leq n}$. Then for large $n$ $M_{n}$ satisfies a recurrence relation $M_{n}=\sum\limits_{j=1}^{s}R_{nj}(\{h_{ab}\})M_{n-j}$ where $s=(l+r)!/l!r!$ is the order of recurrence relation and coefficients $R_{nj}$ satisfy $R_{nj}(\{xh_{ab}\})=x^{j}R_{nj}(\{h_{ab}\})$.
@@
$M_{n}=\sum\limits_{j=1}^{s}R_{nj}(\{h_{ab}\})M_{n-j}.$ (57)
Note that coefficients $c_{i,m}(n-j+1)$ in Eq. (56) satisfy $c_{i,m}(n-j+1)(\{xh_{ab}\})=xc_{i,m}(n-j+1)(\{h_{ab}\})$, the homogeneous property of coefficients in Eq. (57) $R_{nj}(\{xh_{ab}\})=x^{j}R_{nj}(\{h_{ab}\})$ is ensured by the Gaussian elimination. ∎
When the non-Hermitian model is a single-band model with the lattice translational symmetry, namely, we have $h_{ab}=h_{a+1,b+1}$ for any matrix element $h_{ab}$, the coefficients of the recurrence relation $R_{nj}(\{h_{ab}\})$ in Eq. (57) will be independent of $n$: $R_{nj}(\{h_{ab}\})\equiv C_{j}(\{h_{ab}\})$. One can use the following algorithm to obtain the recurrence relation coefficients $R$ based on Theorem 1:
@@
The OBC spectrum of a single-band non-Hermitian Hamiltonian with the translational symmetry of a lattice vector $H=\sum\limits_{ij}c_{i}^{\dagger}h_{ij}c_{j}$ is determined from solutions of the determinant $\det(E-H)$. Denote the determinant of $N$-site lattice model as $D_{N}:=\det(EI_{N}-H_{N})$, then $D_{N}(E)$ is a polynomial of degree $N$ and satisfies the following recurrence relation according to Theorem 1:
@@ Appendix B Recurrence method for one-dimensional periodic nearest-neighbor hopping model
Note that both $\text{Tr}(U)$ and $\det(U)$ are size-independent, so we can apply the recurrence relation method in the main text to obtain the OBC spectrum. For $k=3$ case with identical nearest-neighbor hopping $t_{1}=t_{2}=t_{3}\equiv u_{1}$, $t_{1}^{\prime}=t_{2}^{\prime}=t_{3}^{\prime}\equiv u_{2}$, the coefficients of the recurrence relation in Eq. (74) are: $\text{Tr}(U)=E^{3}-E^{2}(V_{1}+V_{2}+V_{3})+E(-3u_{1}u_{2}+V_{1}V_{2}+V_{1}V_{3}+V_{2}V_{3})+u_{1}u_{2}(V_{1}+V_{2}+V_{3})-V_{1}V_{2}V_{3}$, $\det(U)=u_{1}^{3}u_{2}^{3}$. For a general $k$, we can derive $\text{Tr}(U)$ and $\det(U)$ via the following matrix identity:
$\begin{pmatrix}x_{1}&1\\ y_{1}&0\end{pmatrix}\begin{pmatrix}x_{2}&1\\ y_{2}&0\end{pmatrix}\cdots\begin{pmatrix}x_{n}&1\\ y_{n}&0\end{pmatrix}=\begin{pmatrix}f_{n}(x_{1},\cdots,x_{n},y_{1},\cdots,y_{n})&f_{n-1}(x_{1},\cdots,x_{n-1},y_{1},\cdots,y_{n-1})\\ y_{1}f_{n-1}(x_{2},\cdots,x_{n},y_{2},\cdots,y_{n})&y_{1}f_{n-2}(x_{2},\cdots,x_{n-1},y_{2},\cdots,y_{n-1})\end{pmatrix},$ (75)
where the bivariate polynomials $f_{n}(x_{1},\cdots,x_{n},y_{1},\cdots,y_{n})$ satisfy the following recurrence relation:
$f_{n}(x_{1},\ldots,x_{n},y_{1},\ldots,y_{n})=f_{n-1}(x_{1},\ldots,x_{n-1},y_{1},\ldots,y_{n-1})x_{n}+f_{n-2}(x_{1},\ldots,x_{n-2},y_{1},\ldots,y_{n-2})y_{n}.$ (76)
Let us define $f_{-1}=0,f_{0}=1,f_{i}=f_{i}(x_{1},\ldots,x_{i},y_{1},\ldots,y_{i}),f_{i}^{\prime}=f_{i}(x_{2},\ldots,x_{i+1},y_{2},\ldots,y_{i+1})$, we can readily prove Eq. (76) by induction, since:
$\begin{pmatrix}f_{n-1}&f_{n-2}\\ y_{1}f_{n-2}^{\prime}&y_{1}f_{n-3}^{\prime}\end{pmatrix}\begin{pmatrix}x_{n}&1\\ y_{n}&0\end{pmatrix}=\begin{pmatrix}f_{n}&f_{n-1}\\ y_{1}f_{n-1}^{\prime}&y_{1}f_{n-2}^{\prime}\end{pmatrix}.$ (77)
Applying Eq. (75) to $\text{Tr}(U^{T})$, we have:
$\text{Tr}(U^{T})=f_{k}(E-V_{1},\cdots,E-V_{k},-t_{k}t_{k}^{\prime},-t_{1}t_{1}^{\prime},\cdots,-t_{k-1}t_{k-1}^{\prime})-t_{k}t_{k}^{\prime}f_{k-2}(E-V_{2},\cdots,E-V_{k-1},-t_{1}t_{1}^{\prime},\cdots,-t_{k-2}t_{k-2}^{\prime}).$ (78)
From Eq. (76), one finds:
$f_{k}(E-V_{1},\cdots,E-V_{k},-t_{k}t_{k}^{\prime},-t_{1}t_{1}^{\prime},\cdots,-t_{k-1}t_{k-1}^{\prime})$ $=\Delta_{1,k},$
$f_{k-2}(E-V_{2},\cdots,E-V_{k-1},-t_{1}t_{1}^{\prime},\cdots,-t_{k-2}t_{k-2}^{\prime})$ $=\Delta_{2,k-1},$ (79)
where:
Thus, the recurrence relation in Eq.(74) is written as:
@@
which gives the recurrence relation in the main text.
@@ Appendix D Non-Hermitian Hofstadter butterfly on the triangular lattice
where the magnetic phase factors along the three sides of a triangular plaquette are given by